Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

Out of 2 broken TV sets and 6 fixed TV sets in Joe's Fix-My-

Expert replies
by BTGmoderatorDC » Fri Feb 08, 2019 6:04 am
Timer
00:00
Your answer
Difficulty–

right wrong

Pick an answer to see how members did

Out of 2 broken TV sets and 6 fixed TV sets in Joe's Fix-My-LCD store, 2 sets are selected randomly. What is the probability that at least 1 set is broken?

A. 11/28
B. 13/28
C. 15/28
D. 1/5
E. 1/8

OA B

Source: Economist Gmat
Join the discussion
Source: — Problem Solving |

by Brent@GMATPrepNow » Fri Feb 08, 2019 6:18 am
BTGmoderatorDC wrote:Out of 2 broken TV sets and 6 fixed TV sets in Joe's Fix-My-LCD store, 2 sets are selected randomly. What is the probability that at least 1 set is broken?

A. 11/28
B. 13/28
C. 15/28
D. 1/5
E. 1/8

OA B

Source: Economist Gmat
There are 8 TV's in total
2 are broken
6 are fixed

We want to find P(at least one TV is broken)
When it comes to probability questions involving at least, it's often best to try using the complement.
That is, P(Event A happening) = 1 - P(Event A not happening)
So, here we get: P(getting at least 1 broken TV) = 1 - P(not getting at least 1 broken TV)
What does it mean to not get at least 1 broken TV? It means getting zero broken TVs.

So, we can write: P(getting at least 1 broken TV) = 1 - P(getting zero broken TVs)
1 - P(getting two FIXED TVs)

P(getting two FIXED TVs)
P(getting two FIXED TVs) = P(1st TV fixed and 2nd TV is fixed)
= P(1st TV fixed) x P(2nd TV is fixed)
= 6/8 x 5/7
= 30/56
= 15/28

So, P(getting at least 1 broken TV) = 1 - P(not getting at least 1 broken TV)
= 1 - 15/28
= 13/28

Answer: B

Cheers,
Brent
Brent Hanneson - Creator of GMATPrepNow.com
Image
Join the discussion

by Scott@TargetTestPrep » Wed Feb 13, 2019 6:35 pm
BTGmoderatorDC wrote:Out of 2 broken TV sets and 6 fixed TV sets in Joe's Fix-My-LCD store, 2 sets are selected randomly. What is the probability that at least 1 set is broken?

A. 11/28
B. 13/28
C. 15/28
D. 1/5
E. 1/8

OA B

Source: Economist Gmat
We are given that there are 2 broken TV sets and 6 fixed TV sets and need to determine the probability that when 2 TV sets are selected, at least 1 is broken. Recall that the phrase “at least one” means “one or more.”

We can use the following formula:

1 = P(at least 1 set is broken) + P(no sets are broken)

Thus:

P(at least 1 set is broken) = 1 - P(no sets are broken)

P(no sets are broken) = 6/8 x 5/7 = 3/4 x 5/7 = 15/28

P(at least 1 set is broken) = 1 - 15/28 = 13/28

Answer: B

Scott Woodbury-Stewart
Founder and CEO
[email protected]

Image

See why Target Test Prep is rated 5 out of 5 stars on BEAT the GMAT. Read our reviews

ImageImage
Join the discussion